OpenStudy (anonymous):

1)Determine algebraically whether the function is even, odd, or neither even nor odd. f(x)= x+12/x 2)Find the inverse of the function. f(x) = 5x^3 - 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint for #1 a function is even if f(-x) = f(x) for all x in the domain of f(x) a function is odd if f(-x) = -f(x) for all x in the domain of f(x)

OpenStudy (anonymous):

odd right?

jimthompson5910 (jim_thompson5910):

first off, is f(x) \[\Large f(x) = x+\frac{12}{x}\] or is it \[\Large f(x) = \frac{x+12}{x}\]

OpenStudy (anonymous):

ohh I thought it was odd if all the exponents were odd

OpenStudy (anonymous):

2nd one

jimthompson5910 (jim_thompson5910):

if \[\Large f(x) = \frac{x+12}{x}\] then f(-x) = ???

OpenStudy (anonymous):

sorry, first one

jimthompson5910 (jim_thompson5910):

so it's \[\Large f(x) = x+\frac{12}{x}\]

jimthompson5910 (jim_thompson5910):

right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

if \[\Large f(x) = x+\frac{12}{x}\] then f(-x) = ???

OpenStudy (anonymous):

f(x)

OpenStudy (anonymous):

so even

jimthompson5910 (jim_thompson5910):

tell me what the right side is for f(-x)

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = x+\frac{12}{x}\] \[\Large f(-x) = (-x)+\frac{12}{-x}\] \[\Large f(-x) = ???\]

OpenStudy (anonymous):

negatives cancel out, so it would be even

jimthompson5910 (jim_thompson5910):

how are they cancelling?

OpenStudy (anonymous):

F(-x)=-x?

jimthompson5910 (jim_thompson5910):

how would you simplify \[\Large f(-x) = (-x)+\frac{12}{-x}\]

OpenStudy (anonymous):

Isnt it already simplified? Im confused

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

12/(-x) would turn into -12/x

OpenStudy (anonymous):

ohhh

jimthompson5910 (jim_thompson5910):

so we get \[\Large f(x) = x+\frac{12}{x}\] \[\Large f(-x) = (-x)+\frac{12}{-x}\] \[\Large f(-x) = -x-\frac{12}{x}\]

jimthompson5910 (jim_thompson5910):

So it's clear that f(x) = f(-x) is false

OpenStudy (anonymous):

ohhh I understand now, so its even

jimthompson5910 (jim_thompson5910):

however, we can factor out a negative 1 which leads us to \[\Large f(x) = x+\frac{12}{x}\] \[\Large f(-x) = (-x)+\frac{12}{-x}\] \[\Large f(-x) = -x-\frac{12}{x}\] \[\Large f(-x) = -\left(x+\frac{12}{x}\right)\] \[\Large f(-x) = -f(x)\]

jimthompson5910 (jim_thompson5910):

go back to my definitions I posted above

OpenStudy (anonymous):

By your definition, it's false, correct?

OpenStudy (anonymous):

odd*

jimthompson5910 (jim_thompson5910):

what's false?

OpenStudy (anonymous):

Ohh I didnt know I could factor out the one

OpenStudy (anonymous):

I meant it is odd

jimthompson5910 (jim_thompson5910):

it is odd and hopefully you see how

OpenStudy (anonymous):

yes i got it, thank you so much

jimthompson5910 (jim_thompson5910):

yw